OpenStudy (anonymous):

A shopkeeper buys a certain number of pens. If the selling price of 5 pens is equal to the cost price of 7 pens, find his profit or loss percentage.

OpenStudy (***[isuru]***):

if the selling price of 5 pens = 5x the selling price of one pen = x then the cost of 7 pens also = 5x so the cost of 1 pen = 5x/7 the difference between the selling price and the cost = x - 5x/7 = 2x/7 so the percentage profit = (the difference between the selling price and cost / cost price ) x 100%

OpenStudy (***[isuru]***):